3. If production volume increases from 8000 to 10 000 units, 

a. total costs will increase by 20 per cent. 

b. total costs will increase by 25 per cent. 

c. total variable costs will increase by 25 per cent. 

d. mixed and variable costs will increase by 25 per cent.


4. The direct material cost is £10 000 when 2000 units are produced. What is the direct material cost for 2500 units produced? 

a. £10 000 b. £8000 c. £15 000 d. £12 500 


5. Fixed cost per unit is £9 when 20 000 units are produced and its £6 when 30 000 units are produced. What is the total fixed cost when nothing is produced? 

a. £120 000 b. £270 000 c. £15 d. £180 000 


6. Harry has just received his bachelor's degree and is considering two alternatives. (1) He could obtain an entry-level accounting position paying £30 000 per year. (2) He could obtain his master's degree with one more year of study and work part-time for £8 000 per year. The opportunity cost associated with Harry obtaining his master's degree is 

a. £0. b. £22 000. c. £30 000. d. £38 000.
Answer #1

Solution -

3) if the production volume increases from 8000 units to 10000 units- the variable cost will increase by 25%
for example- let's suppose variable cost per unit is £2

then variable cost for 8000 units =8000*2=£16,000(i)

variable cost for 10000 units =10,000*2=£20,000 (ii)

increase in variable cost( in%)

=[(change in variable cost(ii)-(i))/ variable cost in beginning (when 8000 units were produced )]*100

=[4000/16000]*100=25%

4)answer is d

direct material cost per unit

=direct material cost/number of units

=10000/2000=£5

direct material cost for 2500 units

= direct material cost per unit*number of units produced

=5*2500=£12,500

5)answer is d

always remember total fixed cost remains same irrespective of number of units produced but fixed cost per unit changes every time with change in number of units produced

Total fixed cost =fixed cost per unit *number of units

total fixed cost for 20000units= 9*20000=£1,80,000

Total fixed cost for 30000 units=6*30000=£1,80,000

you can see above that fixed cost per unit is different for every unit but total fixed cost is same

6)answer is c

opportunity cost means the cost of the next best alternative forgone

According to the question , next best alternative was choosing the entry level job which would have given£30000 a year. As this alternative is not chosen, so £30000 is the opportunity cost associated with obtaining masters degree.